# A 60-Addition, Rank-23 Scheme for Exact $\times3$ Matrix Multiplication

### v1

@@ 1 Introduction
Matrix multiplication runs the modern world. For every word that ChatGPT writes, we estimate that $\mathcal{O}(10^{10})$ small matrix products must be computed. [^1] Using the token-level FLOP model $C_{\text{forward}}=2N$ derived by Kaplan et al.[6], a 175 B-parameter model needs $\approx 3.5\times 10^{11}$ fused-multiply-adds per generated token, i.e. $O(10^{10})$ micro-matrix products when mapped to the 32 × 32 kernels used in modern GPU/TPU GEMMs. Due to the extreme volume of matrix multiplication required for most digital workflows, even marginal improvements in fundamental matrix multiplication algorithms can translate to substantial cost and energy savings globally.
In a fast-matrix-multiplication algorithm, multiplicative complexity (often called the rank) is the number of scalar products performed, additive complexity is the total scalar additions or subtractions used to form those products and assemble the result, and arithmetic complexity is simply the sum of the two, i.e. the total scalar operations executed. The quest for ever-smaller arithmetic circuits for matrix multiplication has motivated both theoretical bounds and practical kernels since Strassen’s landmark $2\times 2$ decomposition. In 1969, Strassen [12] found a new algorithm that reduced the ’rank’ (loosely, the number of multiplications) required for computing an arbitrary $\times2$ matrix product from the naive $O(n^{3})$ = 8 to 7, initiating an intensive effort to discover faster matrix multiplication algorithms that continues today. Laderman (1976) [7], discovered a rank-$23$ algorithm for the $3\times 3$ case: the best known upper bound on multiplicative complexity for this problem to this day. With the multiplicative rank seemingly frozen at 23 for almost five decades, later work pivoted to trimming the additive cost of Laderman-type schemes, using graph-theoretic analyses and automated search heuristics [3, 5, 9]. Schwartz and Vaknin subsequently reduced Laderman’s $98$-add form to 61 additions by leveraging an alternative basis and a pebbling heuristic [11].
@@ 2 Neural Discovery Pipeline Overview
At train-time, the networks are initially unconstrained, allowing the weights to explore a dense floating-point search space and reduce a simple mean-squared-error loss until the underlying physics of matrix multiplication has been grokked. Once appropriate accuracy is reached, we activate a ’ternarisation schedule’: a temperature-controlled regularizer over a custom loss function that gradually snaps every weight to the ternary alphabet {-1,0,1} while preserving accuracy as much as possible. Once the network is $\geq$99% ternary, we hard-quantize, read out the network’s integer-valued weight matrices, and verify symbolically that they implement a valid low-rank matrix multiplication scheme. Finally, a Greedy-Potential post-processor from Mårtensson-Wagner reorders the linear parts to minimise redundancies, yielding an exact algorithm.
Our result tightens the best known arithmetic upper bound for exact $3\times 3$ multiplication from $23+61=84$ (Schwartz-Vaknin use a change of basis to achieve this) to $23+60=83$ scalar operations. Unlike Schwartz-Vaknin, our result does not require a change of basis. Unlike multiplicative complexity in the $\times3$ case, the theoretical lower bound on additive complexity is not known (at least, based on our best research efforts), and so this result represents an existence proof of a lower bound.
@@ 6 Discussion
Because multiplications overwhelmingly dictate runtime on modern hardware, trimming two additions from a $\times3$ scheme is unlikely to shift real-world performance; the $2/62\approx 3.2$% reduction in the linear phase is negligible for production BLAS (Basic Linear Algebra Subprograms) kernels, and the extra memory traffic (register shuffles, cache moves, and pointer chasing) tends to overshadow (or even nullify) any arithmetic saving.

### v2

@@ 1 Introduction
Matrix multiplication runs the modern world. For every word that ChatGPT writes, we estimate that $\approx 10^{10}$ small matrix products must be computed. [^1] Using the token-level FLOP model $C_{\text{forward}}=2N$ derived by Kaplan et al.[6], a 175 B-parameter model needs $\approx 3.5\times 10^{11}$ fused-multiply-adds per generated token, i.e. $O(10^{10})$ micro-matrix products when mapped to the 32 $\times$ 32 kernels used in modern GPU/TPU GEMMs. Due to the extreme volume of matrix multiplication required for most digital workflows, even marginal improvements in fundamental matrix multiplication algorithms can translate to substantial cost and energy savings globally.
In a fast-matrix-multiplication algorithm, multiplicative complexity (often called the rank) is the number of scalar products performed, additive complexity is the total scalar additions or subtractions used to form those products and assemble the result, and arithmetic complexity is simply the sum of the two, i.e. the total scalar operations executed. The quest for ever-smaller arithmetic circuits for matrix multiplication has motivated both theoretical bounds and practical kernels since Strassen’s landmark $2\times 2$ decomposition. In 1969, Strassen [12] found a new algorithm that reduced the ’rank’ (loosely, the number of multiplications) required for computing an arbitrary $\times2$ matrix product from the naive $O(n^{3})$ = 8 to 7, initiating an intensive effort to discover faster matrix multiplication algorithms that continues today. Laderman (1976) [7], discovered a rank-$23$ algorithm for the $3\times 3$ case: the best known upper bound on multiplicative complexity for this problem to this day. With the multiplicative rank seemingly frozen at 23 for almost five decades, later work pivoted to trimming the additive cost of Laderman-type schemes, using graph-theoretic analyses and automated search heuristics [3, 5, 9]. Schwartz and Vaknin subsequently reduced a rank-23 scheme to 61 additions by leveraging an alternative basis and a pebbling heuristic [11].
@@ 2 Neural Discovery Pipeline Overview
At train-time, the bilinear networks are initially unconstrained, allowing the weights to explore a dense floating-point search space and reduce a simple mean-squared-error loss until the underlying physics of matrix multiplication has been grokked. We train the network by standard backpropagation: for each minibatch $(A,B)$ we compute $\hat{C}=f_{\theta}(A,B)$ and minimize $\ell(\theta)=\|\hat{C}-AB\|_{F}^{2}+\lambda_{T}\,\mathrm{SmoothMin}_{\mathcal{A}}(\theta;T)$ with AdamW [8], updating $\theta\leftarrow\mathrm{AdamW}(\theta,\nabla_{\theta}\ell)$.
Once appropriate accuracy is reached, we activate a ’ternarisation schedule’: a temperature-controlled regularizer over a custom loss function that gradually snaps every weight to the ternary alphabet {-1,0,1} while preserving accuracy as much as possible; the temperature $T$ is annealed over $R$ steps to push $\theta$ toward $\{-1,0,1\}$. Once the network is $\geq$99% ternary, we hard-quantize, read out the network’s integer-valued weight matrices, and verify symbolically that they implement a valid low-rank matrix multiplication scheme. Finally, a Greedy-Potential post-processor from Mårtensson-Wagner reorders the linear parts to minimise redundancies, yielding an exact algorithm.
Our result tightens the best known arithmetic upper bound for exact $3\times 3$ multiplication from $23+61=84$ (Schwartz-Vaknin use a change of basis to achieve this) to $23+60=83$ scalar operations. Unlike Schwartz-Vaknin, our result does not require a change of basis. Additive optimality in the $3\times 3$ remains open: we are not aware of any nontrivial lower bound pinning the minimum additions for rank-23. Hence, our result is best described as a new constructive upper bound on the arithmetic cost.
@@ 6 Discussion
Trimming two additions from a $\times3$ scheme is unlikely to shift real-world performance without complementary optimisations; the $2/62\approx 3.2$% reduction is negligible for production BLAS (Basic Linear Algebra Subprograms) kernels given that extra memory traffic (register shuffles, cache moves, and pointer chasing) tends to overshadow (or even nullify) any arithmetic saving.